OpenStudy (anonymous):

Poisson distirubution: lambda=2 what is P(Y>=4|Y>=2)??

OpenStudy (anonymous):

P(Y>=4|Y>=2) = P(Y>=2)

OpenStudy (anonymous):

so then i could caluate that as 1-P(Y<2)??

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

They are completely overlapping events.

OpenStudy (anonymous):

Is it given or or?

OpenStudy (anonymous):

yes, but \(Y>4\) is contained in \(Y>2\)

OpenStudy (anonymous):

i could easily be wrong, but i would say it was \[\frac{P(y>4)}{P(y>2)}\]

OpenStudy (anonymous):

yes that is what the book has, ecxpet if >= for both num and denom

OpenStudy (anonymous):

its*

OpenStudy (anonymous):

but how do you know that P(>=4) goes in the numerator?

OpenStudy (anonymous):

\[ \Pr(A \text{ if } B) =\frac{\Pr(A\text{ and }B)}{\Pr(B)} \]

OpenStudy (anonymous):

so its like saying A^B=A if A is a subset of B

OpenStudy (anonymous):

yes, it is \[P(A|B)=\frac{P(A\cap B)}{P(B)}\] but in this case \(A\cap B=A\)

OpenStudy (anonymous):

Oh okay, what if A was not a subset of B?

OpenStudy (anonymous):

then it is what @wio and i both wrote above

OpenStudy (anonymous):

okay, i believe i got it, what about this formula that i found e^((-lambda)*(t))

OpenStudy (anonymous):

since in this case \(A\cap B=A\) it is just \(P(A|B)=\frac{P(A)}{P(B)}\)

OpenStudy (anonymous):

\[P(y\geq 2)=1-P(y=0)-P(y=1)\]

OpenStudy (anonymous):

is that the calculation for the bottom portion?

OpenStudy (anonymous):

\[P(y=0)=e^{-\lambda}\]\[P(y=1)=\lambda e^{-\lambda}\]

OpenStudy (anonymous):

yeah that is the bottom

OpenStudy (anonymous):

alrigh

OpenStudy (anonymous):

i think i got it from here, th

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

:) thanks

OpenStudy (anonymous):

you also need \[P(y=2)=\frac{\lambda^2e^{-\lambda}}{2}\] and \[P(y=3)=\frac{\lambda^3e^{-\lambda}}{6}\]

OpenStudy (anonymous):

general term is \[P(x=k)=\frac{\lambda^ke^{-\lambda}}{k!}\]

OpenStudy (anonymous):

basically to calulate that fraction you are doing: (1-P(y<4)/(1-P(y<2)